OpenStudy (uri):

John is making himself a lunch. He has three different soups to choose from, four kinds of soda, and three different kinds of fruit. How many different lunches can he prepare today? a.28 b.36 c.40 d.48 I need explanation!

OpenStudy (anonymous):

B 3*4*3 Bruh

OpenStudy (uri):

I need explanation.

OpenStudy (uri):

@satellite73 @mathstudent55

OpenStudy (anonymous):

counting principle all the way

OpenStudy (uri):

I totally understand what you just said..

OpenStudy (primeralph):

Just count all the bases, you get what I mean. The answer is 3*4*3.

OpenStudy (anonymous):

if there are \(n\) ways to do one task, and \(m\) ways to do another, then the total number of ways they can be done together is \(n\times m\)

OpenStudy (anonymous):

so if there are 3 ways to do one thing and 4 to do another, total number of ways to do them both is \(3\times 4=12\) this is the COUNTING PRINCIPLE

OpenStudy (uri):

Get it till here,Next?

OpenStudy (anonymous):

which of course extends to more than just two things

OpenStudy (uri):

I got it,Thanks. (:

OpenStudy (anonymous):

if there are \(n_1, n_2, n_3, ...,n_p\) ways to do respectively \(1,2,3,...,p\) tasks, then the number of ways to do them combined is \[n_1\times n_2\times n_3\times ...\times n_p \]

OpenStudy (primeralph):

Shall we go skinny dipping now?